OpenStudy (anonymous):

the density of a solution is 1.18g/ml. find it;s specific gravity

OpenStudy (anonymous):

1.18

OpenStudy (anonymous):

1.18g/ml=1180g/L specific gravity=1180g/L /density of water....=1.18g/L

OpenStudy (anonymous):

specific gravity has no units

OpenStudy (anonymous):

true...sorry